Student in a highschool can choose to study spanish and french as a foriegn language.In a given academic year, 90%of pupils choose spanish and the rest choose french.30%of the student who are learning spanish are boys and 40% of who are studying french also boy. If student was randomly selected from this academic year, What is the probability that it would be





a) a girl?





b) What is the probablity selected boy student given that spanish language study?





c) A boy?





d) what is the probability that a randomly selected student is a boy or somebody who studies french?





e) whay is the probability that randomly selected student is girl who does not study french?

1
Expert's answer
2022-01-24T03:23:45-0500

a)

"P=0.9\cdot0.7+0.1\cdot0.6=0.69"


b)

in class of spanish 30% are boys

so:

"P=0.3"


c)

"P=0.9\cdot0.3+0.1\cdot0.4=0.31"


d)

"P=0.31+0.1-0.1\cdot0.4=0.37"


e)

90%of pupils choose spanish where 70% are girls

so:

"P=0.9\cdot0.7=0.63"